OpenStudy (scarlettfarra2000):

Can someone please show me step by step what to do? Please I'm really having a hard time understanding what it is I'm doing. Simplify the radical expression. Assume that all variables represent positive real numbers.

OpenStudy (scarlettfarra2000):

\[2\sqrt{45z ^{2}a}+10z \sqrt{125a}\]

OpenStudy (mathmale):

Hi, Scarlett, What did you learn from the previous problem you posted here? By now it should be clear that you need to examine the quantities under the radicals and determine the largest perfect square in each radicand. Have you done that yet?

OpenStudy (welshfella):

factor 45 and 125 first so as to simplify the radicals 45 = 5 * 9 125 = 5 * 25 so sqrt125 = 5 sqrt5 and sqrt45 = 3 sqrt5

OpenStudy (mathmale):

What is the largest perfect square factor of 45? What is the same for z^2? What is the same for 125?

OpenStudy (mathmale):

Combine my questions with welshfella's demo and answers.

OpenStudy (scarlettfarra2000):

\[125=5\sqrt{5}\] \[45=3\sqrt{5}\]

OpenStudy (welshfella):

right - now what is the square root of z^2?

OpenStudy (scarlettfarra2000):

2?

OpenStudy (welshfella):

??? does 2 * 2 = z^2

OpenStudy (scarlettfarra2000):

4z

OpenStudy (welshfella):

still wrong 4z * 4z = 16z^2

OpenStudy (welshfella):

square root of z^2 is simply z

OpenStudy (welshfella):

because z * z = z^2 by definition

OpenStudy (welshfella):

so can you simplify the first term 2sqrt ( 45z^2a)

OpenStudy (scarlettfarra2000):

\[2\sqrt{9\times5x ^{2}?}\]

OpenStudy (welshfella):

its 2 * 3 * z * sqrt(5a)

OpenStudy (welshfella):

do you follow that?

OpenStudy (scarlettfarra2000):

Okay back up really quick I'm lost

OpenStudy (scarlettfarra2000):

okay but wouldn't 3*2=6 and 5*5=25 sqrt z?

OpenStudy (welshfella):

Oh whats wrong with me today - forget that 5!!! 2 *3 sqrt5 * z * sqrta

OpenStudy (scarlettfarra2000):

okay I'm starting to see it the problem

OpenStudy (welshfella):

sqrt5 * sqrta = sqrt 5a so we have 6 z sqrt (5a)

OpenStudy (welshfella):

OK?

OpenStudy (scarlettfarra2000):

Yes

OpenStudy (welshfella):

the second term is a little easier can you figure it out?

OpenStudy (scarlettfarra2000):

Maybe

OpenStudy (welshfella):

you already now that sqrt 125 = 5 sqrt5

OpenStudy (scarlettfarra2000):

right

OpenStudy (scarlettfarra2000):

Are we just trying to figure out what's in the square root first? Or do we add the 10 in the problem too?

OpenStudy (welshfella):

10 z * 5 * sqrt 5 * sqrt a

OpenStudy (welshfella):

- simplifying the square root bit first

OpenStudy (welshfella):

check how we did the first term - this is very similar

OpenStudy (scarlettfarra2000):

okay the first one was 3*2 sqrt 5* z * sqrt a in this one the numbers are different it would be 5*2 sqrt 5 * 2 * sqrt a right?

OpenStudy (welshfella):

No not sure what you are thinking there we have 10 *z * 5 * sqrt5 * sqrta first simplify the radical part then the other bits

OpenStudy (scarlettfarra2000):

so do don't simplify 10 when you find the radical (which I'm having a very help time doing)

OpenStudy (welshfella):

sqrt5 * sqrt a = sqrt ?

OpenStudy (welshfella):

its simply sqrt (5a)

OpenStudy (welshfella):

and 10 *z * 5 = what?

OpenStudy (scarlettfarra2000):

50*5 maybe

OpenStudy (welshfella):

No why does the z disappear?

OpenStudy (scarlettfarra2000):

I meant 50*z sorry

OpenStudy (welshfella):

right which is written as 50z

OpenStudy (welshfella):

so we are lef twith 6 z sqrt5 + 50 z sqrt5 - 2 like terms so we just add them for the final answer

OpenStudy (scarlettfarra2000):

\[56\sqrt{5}\]

OpenStudy (welshfella):

Why dont you like the z????

OpenStudy (scarlettfarra2000):

would the z go before 56 or after 5?

OpenStudy (welshfella):

after the 56

OpenStudy (scarlettfarra2000):

ok so like \[56x \sqrt{5}\]

OpenStudy (welshfella):

yea except it z not x

OpenStudy (scarlettfarra2000):

right sorry I'm use to x's not z's

OpenStudy (welshfella):

gotta go bye

OpenStudy (scarlettfarra2000):

Okay thanks for sticking with me I know I asked a lot of stupid questions

OpenStudy (scarlettfarra2000):

@Atrineas